# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v4

@@ A Background
B Proofs of Theorems
@@
Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ is transformed as $\mathbf{Q}(\mathbf{R})$:
$\mathbf{Q}(\mathbf{R})=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger},$
where $\mathbf{D}^{l_{p}}(\mathbf{R})$ and $\mathbf{D}^{l_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l_{p}$ and $l_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
@@
$\mathbf{Q}(\mathbf{R})^{\dagger}=\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}.$
Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
$\mathbf{T}(\mathbf{R})=tr(\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}))$
$=tr(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R}))=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\mathbf{T}.$
Therefore, $\mathbf{T}=tr(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
@@
C Information of Experimental Databases
Figure 2: Illustration of a Hamiltonian matrix. Each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
@@ D Implementation Details
E Implementation Details on the DeepH Benchmark Series
@@
The software environment used is Pytorch 2.0.1 for experiments on the six crystalline databases from the DeepH benchmark series. When combining the proposed TraceGrad method with the DeepH-E3 architecture, the implementation of DeepH-E3 is based on the project  [^3] https://github.com/Xiaoxun-Gong/DeepH-E3 provided by Gong et al. (2023), keeping the architecture and hyperparameter configurations consistent with their setup. In our framework, we use the same number of encoding modules $K$ as DeepH-E3, which is set to $3$. In each encoding module, we apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature, enabling non-linear expressiveness. We set the number of channels for the rotationally SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$; the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $25$ for $MG$ and $BG$, $49$ for $MM$, and $196$ for $BB$, $BT$, and $BS$. It is worth noting that, while $s_{nonlin}(\cdot)$ can be implemented as any differentiable neural network module, we here implement it as a simple fully-connected module. This decision is made to avoid adding significant computational burden to the whole network. Meanwhile, as DeepH-E3 already incorporates complex graph network mechanisms for information aggregation and message-passing, there is no need for $s_{nonlin}(\cdot)$ to be overly complex. Its role is focusing on to filling in the gaps left by the existing equivariant mechanisms in DeepH-E3: to introduce a non-linear mapping mechanism that maintains equivariance, thereby activating and unleashing the expressive power of the overall network architecture through non-linearity. The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix. Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic blocks. Regarding the error metric in the loss function Eq. 7, for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on crystalline structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the six crystalline databases and the searched value is $0.3$. Other hyper-parameters and configurations are the same as DeepH-E3 (Gong et al., 2023): the initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$, and $BS$ databases are set to $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively; the training batch size is set as $1$; the optimizer is chosen as Adam; the scheduler is configured as a slippery slope scheduler.
F Implementation Details on the QH9 Benchmark Series

### v5

@@ A Definition of Concepts
Definition .9.
Direct-Product State. Let $V_{1}$ and $V_{2}$ be two vector spaces, and let $|v_{1}\rangle\in V_{1}$ and $|v_{2}\rangle\in V_{2}$ be arbitrary elements of these spaces. The direct-product state of $|v_{1}\rangle$ and $|v_{2}\rangle$ is defined as the element $|v_{1}\rangle\otimes|v_{2}\rangle$ in the direct-product space $V_{1}\otimes V_{2}$. The direct-product state represents all possible combinations of the elements of $V_{1}$ and $V_{2}$, and forms a new vector space that captures the joint state of two systems. The action of a group on the direct-product state is defined by the direct product of the individual actions on $|v_{1}\rangle$ and $|v_{2}\rangle$, i.e., the effect of the group action on the direct-product space is the direct product of the actions on each individual space:
$g\cdot(|v_{1}\rangle\otimes|v_{2}\rangle)=(g\cdot|v_{1}\rangle)\otimes(g\cdot|v_{2}\rangle),\quad\forall g\in G.$
Definition .10.
Physical Quantity Formed by the Direct-Product of Two Degrees. In quantum mechanics, direct-product spaces are used to describe the combined states of systems with distinct degrees of freedom, such as angular momentum. The combined state captures both the independent action of each degree and their joint transformation under rotations governed by the $\text{SO}(3)$ group.
Let degrees $l_{p}$ and $l_{q}$ represent the angular momentum quantum numbers of two systems. A direct-product state $\mathbf{Q}^{l_{p}\otimes l_{q}}\in\mathbb{R}^{(2l_{p}+1)\times(2l_{q}+1)}$ combines these degrees and transforms under the $\text{SO}(3)$ group according to:
$\mathbf{Q}(\mathbf{R})^{l_{p}\otimes l_{q}}=\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}^{l_{p}\otimes l_{q}}\cdot(\mathbf{D}^{l_{q}}(\mathbf{R}))^{\dagger},$
where $\mathbf{D}^{l_{p}}(\mathbf{R})\in\mathbb{R}^{(2l_{p}+1)\times(2l_{p}+1)}$ and $\mathbf{D}^{l_{q}}(\mathbf{R})\in\mathbb{R}^{(2l_{q}+1)\times(2l_{q}+1)}$ are the Wigner-D matrices for the degrees $l_{p}$ and $l_{q}$, respectively, and $\dagger$ denotes the conjugate transpose, ensuring unitary transformations.
Definition .11.
Direct-Sum State. Let $V_{1}$ and $V_{2}$ be two vector spaces, and let $|v_{1}\rangle\in V_{1}$ and $|v_{2}\rangle\in V_{2}$ be arbitrary elements of these spaces. The direct-sum state of $|v_{1}\rangle$ and $|v_{2}\rangle$ is defined as the element $|v_{1}\rangle\oplus|v_{2}\rangle$ in the direct-sum space $V_{1}\oplus V_{2}$.
The direct-sum space $V_{1}\oplus V_{2}$ consists of ordered pairs of elements, where each element is drawn from one of the original spaces. The operations of vector addition and scalar multiplication in $V_{1}\oplus V_{2}$ are defined component-wise:
$(a_{1},a_{2})+(b_{1},b_{2})=(a_{1}+b_{1},a_{2}+b_{2}),\quad c\cdot(a_{1},a_{2})=(c\cdot a_{1},c\cdot a_{2}),$
for $(a_{1},a_{2}),(b_{1},b_{2})\in V_{1}\oplus V_{2}$ and $c\in\mathbb{F}$, where $\mathbb{F}$ is a field, typically either the real numbers $\mathbb{R}$ or the complex numbers $\mathbb{C}$.
The direct-sum state $|v_{1}\rangle\oplus|v_{2}\rangle$ represents a combination where the components remain in their respective vector spaces and do not interact with each other. The action of a group $G$ on the direct-sum state is defined by its independent action on each component:
$g\cdot(|v_{1}\rangle\oplus|v_{2}\rangle)=(g\cdot|v_{1}\rangle)\oplus(g\cdot|v_{2}\rangle),\quad\forall g\in G.$
Definition .12.
Clebsch-Gordan Decomposition for $\text{SO}(3)$ Group.
In quantum mechanics, the Clebsch-Gordan decomposition is used to describe the coupling and decoupling of two angular momentum systems. For two angular momentum quantum numbers $l_{p}$ and $l_{q}$, their combined state can be decomposed into series components in the direct-sum state. Taking $\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Definition .10 as an example, it can be decomposed as follows:
$CGDecomp(\mathbf{Q}^{l_{p}\otimes l_{q}})=\bigoplus_{l=|l_{p}-l_{q}|}^{l_{p}+l_{q}}\mathbf{q}^{l},\,\,\,q_{m}^{l}=\sum_{m_{p},m_{q}}C_{m,m_{p},m_{q}}^{l,l_{p},l_{q}}Q^{l_{p}\otimes l_{q}}_{m_{p},m_{q}}$
where $CGDecomp(\cdot)$ denotes the Clebsch-Gordan decomposition operator, $Q^{l_{p}\otimes l_{q}}_{m_{p},m_{q}}$ are the elements of the quantity $\mathbf{Q}^{l_{p}\otimes l_{q}}$ in the direct-product state, $C_{m,m_{p},m_{q}}^{l,l_{p},l_{q}}$ are the Clebsch-Gordan coefficients, and $\mathbf{q}^{l}$ denotes the decomposed results, with $q_{m}^{l}$ representing the component under quantum numbers $l$ and $m$.
Figure 2: Illustration of atomic pairwise Hamiltonian matrices partitioned from the complete Hamiltonian matrix of the whole system. Each Hamiltonian matrix contains multiple basic blocks, denoted as $\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}$.
B Application Task Description: Electronic-structure Hamiltonian Calculation
Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has become a cornerstone of modern electronic structure theory, playing a pivotal role in condensed matter physics, quantum chemistry, and materials science. Introduced in the 1960s through the foundational work of Hohenberg, Kohn, and Sham, DFT provides a framework for studying many-electron systems by replacing the computationally expensive many-body wavefunction with the electron density $\rho(\mathbf{r})$ as the fundamental variable. This reformulation significantly reduces computational complexity while preserving essential quantum mechanical effects, enabling researchers to investigate systems of practical interest with manageable computational resources. Over the decades, DFT has proven to be instrumental in calculating orbital energies and band structures, optimizing structural geometries, and exploring a wide range of material properties, underscoring its versatility and importance across diverse scientific disciplines.
At the heart of DFT lies the Kohn-Sham equation (Kohn & Sham, 1965), which simplifies the many-body problem into a set of single-particle equations. These equations are expressed as:
$\hat{H}\psi_{i}(\mathbf{r})=\epsilon_{i}\psi_{i}(\mathbf{r}),\quad\text{subject to}\quad\hat{H}=-\frac{\hbar^{2}}{2m}\nabla^{2}+V_{\text{ext}}(\mathbf{r})+V_{\text{HXC}}[\rho](\mathbf{r})$ (5)
where $\hat{H}$ is the Hamiltonian operator, the Hartree-exchange-correlation potential $V_{\text{HXC}}[\rho](\mathbf{r})=V_{\text{H}}[\rho](\mathbf{r})+V_{\text{XC}}[\rho](\mathbf{r})$ is a functional of the electron density $\rho(\mathbf{r})$. The electron density is obtained from the Kohn-Sham orbitals $\psi_{i}(\mathbf{r})$ as:
$\rho(\mathbf{r})=\sum_{i=1}^{M}|\psi_{i}(\mathbf{r})|^{2},$
where $M$ represents the total number of electrons.
Atomic orbitals offer a computationally efficient basis for electronic structure calculations because they require fewer basis functions compared to other types of basis sets (Lin et al., 2023). Additionally, their inherently localized nature makes them particularly advantageous for large systems. When expressed in an atomic orbital basis set, the Kohn-Sham equations can be formulated into as a generalized eigenvalue problem:
$\mathbf{H}\mathbf{C}=\mathbf{\epsilon}\mathbf{S}\mathbf{C},$ (6)
where $\mathbf{H}$ is the Hamiltonian matrix incorporating the contributions from $V_{\text{ext}}$ and $V_{\text{HXC}}$, $\mathbf{S}$ is the overlap matrix, $\mathbf{C}$ contains the orbital coefficients, and $\mathbf{\epsilon}$ represents the eigenvalues of the system.
The atomic orbitals used as the basis functions typically take the form:
$\phi_{\mu}(\mathbf{r})=R_{\mu}(r)Y_{l_{\mu}m_{\mu}}(\theta,\phi),$
where $R_{\mu}(r)$ is the radial part of the wavefunction, and $Y_{l_{\mu}m_{\mu}}(\theta,\phi)$ are the spherical harmonics describing the angular dependence. These orbitals are localized around their respective atomic centers, decay rapidly as $r$ increases, and are truncated to zero beyond a cutoff radius $r_{c}$.
The matrix elements of the Hamiltonian and overlap matrices, $H_{\mu\nu}$ and $S_{\mu\nu}$, are defined as:
[…]
The localized nature of atomic orbitals ensures that matrix elements are non-zero only when the orbitals $\phi_{\mu}$ and $\phi_{\nu}$ overlap according to their cutoff radius. Beyond this range, the corresponding matrix elements are treated as zero, leading to sparsity in the Hamiltonian and overlap matrices. This sparsity significantly reduces computational effort and memory usage, especially for large systems.
Solving $H_{\mu\nu}$ involves an iterative self-consistent process: starting with an initial guess for $\rho^{(1)}(\mathbf{r})$, the potentials $V_{\text{HXC}}[\rho](\mathbf{r})$ are updated, and the equations are solved repeatedly until $\rho(\mathbf{r})$ converges. This process can be illustrated as $\rho^{(1)}(\mathbf{r})\rightarrow V^{(1)}_{\text{HXC}}[\rho](\mathbf{r})\rightarrow H^{(1)}_{\mu\nu}\rightarrow\psi^{(1)}_{\mu}(\mathbf{r})\rightarrow\rho^{(2)}(\mathbf{r})\rightarrow\ldots\rightarrow\rho^{(T)}(\mathbf{r})$, until the charge density $\rho(\mathbf{r})$ converges. At that point, the Hamiltonian matrix is outputted by $\rho^{(T)}(\mathbf{r})\rightarrow H^{(T)}_{\mu\nu}$, from which downstream physical properties such as orbital energies and band structures are induced, determining the electronic, magnetic, and transport characteristics of the electron system.
Despite the remarkable success of Kohn-Sham DFT in advancing fields such as materials science, energy, and biomedicine over recent decades (Nagy, 1998; Jones, 2015), the challenge of high computational complexity remains unresolved. The main computational bottleneck lies in iterative solving the Kohn-Sham equation, especially in the matrix diagonalization of Eq. (6), which has a complexity of $\mathcal{O}(N^{3})$, where $N$ is the number of atoms in the system. To address this challenge, recent approaches (Li et al., 2022; Yu et al., 2023b; Gong et al., 2023) have applied the deep graph learning paradigm to predict the self-consistent Hamiltonian. These methods use the Hamiltonian matrix $\mathbf{H}^{(T)}$, calculated from traditional DFT methods, as labels. This matrix can be partitioned into a series of atomic pairwise Hamiltonian matrices $\{\mathbf{H}^{ij}\mid j\in\Omega(i)\}$, as shown in Fig. 2, where $i$ and $j$ represent two atoms in the system. Assuming that the angular momentum quantum numbers of the orbitals of atom $i$ are $l^{i}_{p}\in\{l^{i}_{1},l^{i}_{2},...,l^{i}_{P}\}$, and those of atom $j$ are $l^{j}_{q}\in\{l^{j}_{1},l^{j}_{2},...,l^{j}_{Q}\}$, the basic block of $\mathbf{H}^{ij}$ can be expressed as ${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}$, which is formed by the direct product between the $l^{i}_{p}$ and $l^{j}_{q}$.
These methods train an efficient graph neural network to directly predict each $\mathbf{H}^{ij}$ from the 3D structure of the atomic system, thereby circumventing the extremely time-consuming self-consistent iterations. During the inference phase, these methods successfully reduced the computational complexity of calculating the Hamiltonian matrices to $\mathcal{O}(N)$, while showing good potential in generalizing to larger atomic systems, even though the training set, constrained by the expensive DFT labels, only includes smaller systems. Once the Hamiltonian matrices are obtained, many downstream physical properties can be efficiently calculated with $\mathcal{O}(N)$ complexity. A detailed analysis of the computational complexity is provided in Appendix I. This paper aims to tackle the challenge of predicting electronic Hamiltonians with high accuracy and reliability, while rigorously maintaining SO(3)-equivariance.
C Formulation of Hamiltonian Prediction as an SO(3)-Equivariant Non-linear Representation Learning Problem
As shown in Fig. 2, a basic block of Hamiltonian can be denoted as ${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}$, which is formed by the direct product between degrees $l^{i}_{p}$ and $l^{j}_{q}$. It obeys the following SO(3)-equivariant law:
${\bf{H}}({\bf{R}})^{l^{i}_{p}\otimes l^{j}_{q}}={\bf{D}}^{l^{i}_{p}}({\bf{R}})\cdot{\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}\cdot{({\bf{D}}^{l^{j}_{q}}({\bf{R}}))^{\dagger}}$ (7)
where $\dagger$ denotes the conjugate transpose operation. ${\bf{R}}\in\mathbb{R}^{3\times 3}$ is the rotational matrix, ${\bf{D}}^{l^{i}_{p}}({\bf{R}})\in\mathbb{R}^{(2l^{i}_{p}+1)\times(2l^{i}_{p}+1)}$ and ${\bf{D}}^{l^{j}_{q}}({\bf{R}})\in\mathbb{R}^{(2l^{j}_{q}+1)\times(2l^{j}_{q}+1)}$ are the Wigner-D matrices of degrees $l^{i}_{p}$ and $l^{j}_{q}$, respectively; ${\bf{H}}({\bf{R}})^{l^{i}_{p}\otimes l^{j}_{q}}\in\mathbb{R}^{(2l^{i}_{p}+1)\times(2l^{j}_{q}+1)}$ denotes the transformed results of ${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}\in\mathbb{R}^{(2l^{i}_{p}+1)\times(2l^{j}_{q}+1)}$ through the rotational operation by ${\bf{R}}$.
${\bf{H}}^{l^{i}_{p}\otimes l^{j}_{q}}$ in the direct-product state can be further decomposed into a series of direct-sum state components, i.e., ${\bf{h}}^{l}(|l^{i}_{p}-l^{j}_{q}|\leq l\leq l^{i}_{p}+l^{j}_{q})$, which follows SO(3)-equivariant law mathematically equivalent to Eq. (7) but with a simpler form:
${\bf{h}}({\bf{R}})^{l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{h}}^{l},|l^{i}_{p}-l^{j}_{q}|\leq l\leq l^{i}_{p}+l^{j}_{q}$ (8)
where ${\bf{h}}^{l}\in\mathbb{R}^{2l+1}$ and ${\bf{h}}({\bf{R}})^{l}\in\mathbb{R}^{2l+1}$ respectively denote the components with degree $l$ before and after the rotational operation by ${\bf{R}}$.
For ease of processing, the internal representations of SO(3)-equivariant neural networks (Gong et al., 2023; Liao & Smidt, 2023) are typically in the direct-sum form. To obey the equivariant law of Eq. (8) for regressing Hamiltonian, these hidden representations must also satisfy the same form of equivariance:
${\bf{f}}({\bf{R}})^{(k)l}={\bf{D}}^{l}({\bf{R}})\cdot{\bf{f}}^{(k)l}$ (9)
where ${\bf{f}}^{(k)l}\in\mathbb{R}^{2l+1}$ and ${\bf{f}}({\bf{R}})^{(k)l}\in\mathbb{R}^{2l+1}$ respectively denote one channel of hidden representations with degree $l$ before and after the rotational operation by ${\bf{R}}$, at the $k$ th hidden layer.
Due to the intrinsic complexity and non-linearity of Hamiltonians, the regression neural networks are supposed to equip with powerful non-linear mappings as feature encoders to fully capture the intrinsic patterns of the Hamiltonians, which is crucial for precise and generalizable prediction performance. Meanwhile, the non-linear mappings, denoted as $g_{nonlin}(\cdot)$, must also preserve SO(3)-equivariance, which is expressed as:
However, directly implementing $g_{nonlin}(\cdot)$ as neural network module with non-linear activation functions, such as $Sigmoid$, $Softmax$ and $SiLU$, may result in the destruction of strict equivariance. How to make $g_{nonlin}(\cdot)$ both theoretically SO(3)-equivariant and capable of powerful non-linear expressiveness, and effectively apply it to the prediction of SO(3)-equivariant complex physical quantities, i.e., electronic-structure Hamiltonians in this context, is the core problem this paper aims at solving.
D Proofs of Theorems
@@
Under an SO(3) rotation represented by the rotational matrix $\mathbf{R}$, $\mathbf{H}=\mathbf{H}^{l^{i}_{p}\otimes l^{j}_{q}}$ is transformed as $\mathbf{H}(\mathbf{R})$:
$\mathbf{H}(\mathbf{R})=\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{H}\cdot\mathbf{D}^{l^{j}_{q}}(\mathbf{R})^{\dagger},$
where $\mathbf{D}^{l^{i}_{p}}(\mathbf{R})$ and $\mathbf{D}^{l^{j}_{q}}(\mathbf{R})$ are the Wigner-D matrices for the degrees of $l^{i}_{p}$ and $l^{j}_{q}$, respectively, corresponding to the rotation $\mathbf{R}$.
@@
$\mathbf{H}(\mathbf{R})^{\dagger}=\mathbf{D}^{l^{j}_{q}}(\mathbf{R})\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}.$
Using the cyclic property of the trace, which states that the trace of a product of matrices remains unchanged under cyclic permutations (i.e., $tr(ABC)=tr(BCA)=tr(CAB)$), and combining the properties that $\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}=\mathbf{I}$ and $\mathbf{D}^{l^{j}_{q}}(\mathbf{R})\cdot\mathbf{D}^{l^{j}_{q}}(\mathbf{R})^{\dagger}=\mathbf{I}$, we can rearrange the terms inside the trace as follows:
$\mathbf{T}(\mathbf{R})=tr(\mathbf{H}(\mathbf{R})\cdot\mathbf{H}(\mathbf{R})^{\dagger})=tr((\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{H}\cdot\mathbf{D}^{l^{j}_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l^{j}_{q}}(\mathbf{R})\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}))$
$=tr(\mathbf{D}^{l^{i}_{p}}(\mathbf{R})\cdot\mathbf{H}\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger})=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R})^{\dagger}\cdot\mathbf{D}^{l^{i}_{p}}(\mathbf{R}))=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})=\mathbf{T}.$
Therefore, $\mathbf{T}=tr(\mathbf{H}\cdot\mathbf{H}^{\dagger})$ is invariant under SO(3) transformations, its SO(3)-invariance is proved. ∎
@@
E Information of Experimental Databases
@@ F Implementation Details
F.1 Implementation Details on the DeepH Benchmark Series
@@
The software environment used is Pytorch 2.0.1 for experiments on the six crystalline databases from the DeepH benchmark series. When combining the proposed TraceGrad method with the DeepH-E3 architecture, the implementation of DeepH-E3 is based on the project  [^2] https://github.com/Xiaoxun-Gong/DeepH-E3 provided by Gong et al. (2023), keeping the architecture and model hyperparameters consistent with their setup. In our framework, we use the same number of encoding modules $K$ as DeepH-E3, which is set to $3$. In each encoding module, we apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature, enabling non-linear expressiveness. We set the number of channels for the SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$; the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $25$ for $MG$ and $BG$, $49$ for $MM$, and $196$ for $BB$, $BT$, and $BS$. It is worth noting that, while $s_{nonlin}(\cdot)$ can be implemented as any differentiable neural network module, we here implement it as a simple fully-connected module. This decision is made to avoid adding significant computational burden to the whole network. Meanwhile, as DeepH-E3 already incorporates complex graph network mechanisms for information aggregation and message-passing, there is no need for $s_{nonlin}(\cdot)$ to be overly complex. Its role is focusing on to filling in the gaps left by the existing equivariant mechanisms in DeepH-E3: to introduce a non-linear mapping mechanism that maintains equivariance, thereby activating and unleashing the expressive power of the overall network architecture through non-linearity. The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for an atomic pairwise Hamiltonian matrix. Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic matrix blocks. Regarding the error metric in the loss function Eq. (3), for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on crystalline structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the six crystalline databases and the searched value is $0.3$. To ensure the convergence of the TraceGrad method, we set the maximum training epochs to $5,000$. Other hyper-parameters and configurations are the same as DeepH-E3 (Gong et al., 2023): the initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$, and $BS$ databases are set to $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively; the training batch size is set as $1$; the optimizer is chosen as Adam; the scheduler is configured as a slippery slope scheduler.
F.2 Implementation Details on the QH9 Benchmark Series